OpenStudy (anonymous):

An article in a shop marked at $240 was sold for $192. What was the selling price?

OpenStudy (anonymous):

selling price? wasn't that 192 $ o.O

OpenStudy (anonymous):

I mean what is the discount per cent

OpenStudy (anonymous):

ahh.. that makes sense ;) \[\LARGE 240(1-x\%)=192\] \[\LARGE 240(1-\frac{x}{100 })=192\] \[\LARGE 240-\frac{240x}{100 }=192\] \[\LARGE 240-\frac{24x}{10 }=192\] multiply both sides by 10 \[\LARGE 2400-24x=1920\] \[\LARGE -24x=-480 \] can you solve it from here? :)

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

20

OpenStudy (anonymous):

%%%

OpenStudy (anonymous):

why wtf?? what's there to get suprised of ? o.O

OpenStudy (anonymous):

Original price: $240 Sale price: $192 \[240-192=48\] You saved $48 on the product (your discount). Find the percentage of $48 compared to the original price, $240. \[\frac{48}{240}*100=20\] So you received a 20% discount.